A firm has marginal costs given by MC=10+Q and average variable costs AVC= 10+Q/2 If fixed costs are $5000 and the market price is $100 find firm's maximum profit. Will the firm continue to operate in the short run? Explain.
Answer #1

Answer

The firm produces at MC=P if P>AVC

equating MC=P

10+Q=100

Q=90

AVC=10+Q/2=10+90/2=55

P>AVC so the firm operates

TC=Fied cost +AVC*Q=5000+55*90=9950

TR=P*Q=100*90=9000

Profit=TR-TC=9000-9950=-950

the firm makes a loss of $950

=====

The firm continue to operate as if the firm shutdown the loss is equal to fixed cost which is $5000 so the firm operate in the short run to reduce loss to $950.